# Graded Betti numbers of the Jacobian algebra of surfaces in $\mathbb{P}^{3}$

### v1

@@ Abstract.
We compute an explicit closed formula for the Hilbert polynomial of the Jacobian algebra $M(f)$ of a reduced surface $X:f=0$ in $\mathbb{P}^{3}$ in terms of the graded Betti numbers of the algebra $M(f)$. When $X$ has only isolated singularities, a result by A. du Plessis and C. T. C. Wall yields new necessary condition for a set of positive integers to be the graded Betti numbers of the Jacobian algebra of such a surface. The comparison with the plane curve case is discussed in detail and additional information is given in the case of nodal surfaces.
@@ 1. Introduction
One has the following result, see [17, Theorem 5.3].
@@
The existence of curves $C$ as above is shown in [12, Example 4.5] and a complete characterization of them is given in [12, Theorem 3.5 (1)].
@@
This is a special case of Chebyshev hypersurfaces, which are classical examples of nodal hypersurfaces with many singularities. They were introduced by Chmutov to construct complex projective hypersurfaces with a large number of nodes, see [3] Vol. 2, p. 419, [4] as well as [10, Corollary 3.2, (iii)]. The minimal resolution of the Jacobian algebra is given by
@@ 4. Comparison to the curve case
which corresponds to the first equality in Theorem 1.1 (1) above. It follows from [20, Lemma 1.1] that
@@
Using [20, Formula (13)], one obtains the relation
@@
Note that if the surface $X$ is tame with respect to the pair $(\rho_{1},\rho_{2})$ as in [14, Definition 1.2], then the first order Jacobian syzygies $\rho_{1},\rho_{2}$ and $\rho_{3}$ are linearly independent over the field $K$ for any new additional syzygy $\rho_{3}$. However, unlike the curve case, Example 5.5 shows that in general one may have $t(X)<0$.
@@
It is known that for such a curve one has $\epsilon_{j}=1$ for all $j=1,\ldots,q^{\prime}$ and $d_{1}^{\prime}=\ldots=d_{p^{\prime}}$, see [13, Theorem 3.1]. It is interesting to note that the Cayley and the Kummer surfaces considered in Examples 3.4 and 3.5 enjoy similar properties for their graded Betti numbers, that is the $d_{i}$’s are all equal, and $\alpha_{j}=\beta_{k}=1$ for $j=1,\ldots,d-3$ and $k=1,\ldots,r$. It would be nice to have a theoretical explanation for this fact.
@@
follows from [8, Equation (4.5) and Corollary 4.3]. The inequality
@@
(ii) A similar result to Proposition 4.4 may be stated in the more general situation when the surface $X$ has only isolated weighted homogeneous singularities, by using [9, Theorem 9], which extends the results in [8] to this setting.

### v2

@@ Abstract.
We compute an explicit closed formula for the Hilbert polynomial of the Jacobian algebra $M(f)$ of a reduced surface $X:f=0$ in $\mathbb{P}^{3}$ in terms of the graded Betti numbers of the algebra $M(f)$. When $X$ has only isolated singularities, a result by A. du Plessis and C. T. C. Wall yields new necessary condition for a set of positive integers to be the graded Betti numbers of the Jacobian algebra of such a surface. The comparison with the plane curve case is discussed in detail and additional information is given in the case of nodal surfaces. In the final section we construct four natural Jacobian syzygies for surfaces $X$ coming from pencils of surfaces.
@@ 1. Introduction
In the final section we construct four natural Jacobian syzygies for surfaces $X$ coming from pencils of surfaces, similar to the case of plane curves described in [8, Section 4]. This construction allows us to construct surfaces $X$ having the type t(X)<Nt(X)<-N, where $N$ is an arbitrary large positive number, see Example 6.2. We would like to thank Piotr Pokora for bringing this example to our attention.
@@
One has the following result, see [18, Theorem 5.3].
@@
The existence of curves $C$ as above is shown in [13, Example 4.5] and a complete characterization of them is given in [13, Theorem 3.5 (1)].
@@
This is a special case of Chebyshev hypersurfaces, which are classical examples of nodal hypersurfaces with many singularities. They were introduced by Chmutov to construct complex projective hypersurfaces with a large number of nodes, see [3] Vol. 2, p. 419, [4] as well as [11, Corollary 3.2, (iii)]. The minimal resolution of the Jacobian algebra is given by
@@ 4. Comparison to the curve case
which corresponds to the first equality in Theorem 1.1 (1) above. It follows from [21, Lemma 1.1] that
@@
Using [21, Formula (13)], one obtains the relation
@@
Note that if the surface $X$ is tame with respect to the pair $(\rho_{1},\rho_{2})$ as in [15, Definition 1.2], then the first order Jacobian syzygies $\rho_{1},\rho_{2}$ and $\rho_{3}$ are linearly independent over the field $K$ for any new additional syzygy $\rho_{3}$. However, unlike the curve case, Example 5.5 shows that in general one may have $t(X)<0$.
@@
It is known that for such a curve one has $\epsilon_{j}=1$ for all $j=1,\ldots,q^{\prime}$ and $d_{1}^{\prime}=\ldots=d_{p^{\prime}}$, see [14, Theorem 3.1]. It is interesting to note that the Cayley and the Kummer surfaces considered in Examples 3.4 and 3.5 enjoy similar properties for their graded Betti numbers, that is the $d_{i}$’s are all equal, and $\alpha_{j}=\beta_{k}=1$ for $j=1,\ldots,d-3$ and $k=1,\ldots,r$. It would be nice to have a theoretical explanation for this fact.
@@
follows from [9, Equation (4.5) and Corollary 4.3]. The inequality
@@
(ii) A similar result to Proposition 4.4 may be stated in the more general situation when the surface $X$ has only isolated weighted homogeneous singularities, by using [10, Theorem 9], which extends the results in [9] to this setting.
@@ 5. Examples
6. On the Jacobian syzygies of surfaces coming from pencils of surfaces in $\mathbb{P}^{3}$
Let $g,h\in S$ be two homogeneous polynomials of degree $k\geq 2$ and consider the pencil of surfaces in $\mathbb{P}^{3}$ given by
$\mathcal{P}:\alpha g+\beta h=0,$
where $\alpha,\beta\in\mathbb{C}$. Consider a surface
$X:f=q_{1}\ldots q_{m}=0,$
for some integer $m\geq 2$, where $q_{j}=\alpha_{j}g+\beta_{j}h$ are reduced members of the pencil $\mathcal{P}$ considered above. Then $X$ is a reduced surface of degree $d=km$, which has four natural Jacobian syzygies of degree 2k22k-2 that we describe now. Recall that a syzygy
(6.1) $\rho=(a^{x},a^{y},a^{z},a^{t})\text{ corresponding to }a^{x}f_{x}+a^{y}f_{y}+a^{z}f_{z}+a^{t}f_{t}=0,$
with $a^{x},a^{y},a^{z},a^{t}\in S$ may be identified to a differential 3-form
(6.2) $\omega(\rho)=\ a^{x}\operatorname{d}y\wedge\operatorname{d}z\wedge\operatorname{d}t-a^{y}\operatorname{d}x\wedge\operatorname{d}z\wedge\operatorname{d}t+a^{z}\operatorname{d}x\wedge\operatorname{d}y\wedge\operatorname{d}t-a^{t}\operatorname{d}x\wedge\operatorname{d}y\wedge\operatorname{d}z,$
satisfying
(6.3) $\operatorname{d}f\wedge\omega(\rho)=0$
with $\operatorname{d}f=f_{x}\operatorname{d}x+f_{y}\operatorname{d}y+f_{z}\operatorname{d}z+f_{t}\operatorname{d}t$ the differential of $f$. The differential 2-form
(6.4) $\omega(\mathcal{P})=\operatorname{d}g\wedge\operatorname{d}h$
clearly satisfies
(6.5) $\operatorname{d}f\wedge\omega(\mathcal{P})=0,$
a key fact that was used with a similar purpose in the plane curve case in [8]. It follows that the following four 3-forms
(6.6) $\omega^{x}=\operatorname{d}x\wedge\omega(\mathcal{P}),\ \omega^{y}=\operatorname{d}y\wedge\omega(\mathcal{P}),\ \omega^{z}=\operatorname{d}z\wedge\omega(\mathcal{P})\text{ and }\omega^{t}=\operatorname{d}t\wedge\omega(\mathcal{P})$
yield four Jacobian syzygies for the surface $X$. We have the following result, showing that these syzygies are in general linearly independent.
Theorem 6.1.
Assume that the base locus
$B=\{g=h=0\}$
of the pencil $\mathcal{P}$ with $k=\deg g=\deg h\geq 2$ is a reduced curve. If a linear dependence relation
$c^{x}\omega^{x}+c^{y}\omega^{y}+c^{z}\omega^{z}+c^{t}\omega^{t}=0$
with $c^{x},c^{y},c^{z},c^{t}\in\mathbb{C}$ holds, then the curve $B$ is contained in the plane
$H:\ell=c^{x}x+c^{y}y+c^{z}z+c^{t}t=0.$
In particular, the differential forms $\omega^{x},\omega^{y},\omega^{z}$ and $\omega^{t}$ span a vector subspace of dimension $3$ or $4$.
Proof.
First note that the linear dependence relation is equivalent to the equality
(6.7) $\operatorname{d}\ell\wedge\omega(\mathcal{P})=0,$
with $\operatorname{d}\ell$ the differential of $\ell$. Let $B_{1}$ be an irreducible component of the base locus $B$ and let $s$ be a smooth point on $B_{1}$. It follows that the differentials $\operatorname{d}_{s}g$ and $\operatorname{d}_{s}h$ are linearly independent linear forms. The condition (6.7) is equivalent to saying that $\ell$, identified to $\operatorname{d}_{s}\ell$, is a linear combination of $\operatorname{d}_{s}g$ and $\operatorname{d}_{s}h$. Since $\operatorname{d}_{s}g(s)=\operatorname{d}_{s}h(s)=0$, it follows that $\ell(s)=0$. Since $B_{1}$ is a reduced curve, the smooth points are dense in $B_{1}$, hence $B_{1}$ is contained in the hyperplane $H$. This proof applies to any irreducible component of $B$, hence $B\subset H$.
If $k>1$ then $B$ is a complete intersection of degree $k^{2}>1$, hence $B$ cannot be a line. But a reduced curve contained in a line is a line. It follows that there cannot be two independent linear relations involving the 4 differential forms, since otherwise $B\subset H_{1}\cap H_{2}$, where $H_{1}$ and $H_{2}$ are two distinct planes in $\mathbb{P}^{3}$. ∎
The following example shows that the four Jacobian syzygies constructed in Theorem 6.1 may not be primitive, that is they can be the multiples of lower degree Jacobian syzygies.
Example 6.2.
Consider the surface of degree $d=3m$ given by
$X:f=(x^{3}-yzt)^{m}+(t^{3}-xyz)^{m}=0,$
for any $m\geq 2$, that extends Examples 5.4 and 5.5 above. Then $X$ is a union of $m\geq 2$ cubic surfaces in $\mathbb{P}^{3}$ that are reduced members of a pencil $\mathcal{P}$ as above with
$g=x^{3}-yzt\text{ and }h=t^{3}-xyz.$
A direct computation shows that
$\omega(\mathcal{P})=\operatorname{d}g\wedge\operatorname{d}h=-z(3x^{3}+yzt)\operatorname{d}x\wedge\operatorname{d}y-y(3x^{3}+yzt)\operatorname{d}x\wedge\operatorname{d}z$-
$(y^{2}z^{2}-9x^{2}t^{2})\operatorname{d}x\wedge\operatorname{d}t-z(3t^{3}+xyz)\operatorname{d}y\wedge\operatorname{d}t-y(3t^{3}+xyz)\operatorname{d}t\wedge\operatorname{d}t.$
It follows that
$\omega^{x}=\operatorname{d}x\wedge\omega(\mathcal{P})=-(3t^{3}+xyz)(z\operatorname{d}x\wedge\operatorname{d}y\wedge\operatorname{d}t+y\operatorname{d}x\wedge\operatorname{d}z\wedge\operatorname{d}t).$
Using the formulas (6.1) and (6.2), we see that the Jacobian syzygy $\rho^{x}$ corresponding to the differential form $\omega^{x}$ is
$\rho^{x}=(3t^{3}+xyz)\rho_{1}$
where
$\rho_{1}=(0,y,-z,0)$
is the lowest degree syzygy of $f$. A similar computation shows that the Jacobian syzygy $\rho^{t}$ corresponding to the differential form $\omega^{t}$ is
$\rho^{x}=(3x^{3}+yzt)\rho_{1}.$
Moreover, the Jacobian syzygies $\rho^{y}$ and $\rho^{z}$ corresponding to the differential form $\omega^{y}$ and respectively $\omega^{z}$ are given by the following formulas
$\rho^{y}=(-y(3t^{3}+xyz),0,y^{2}z^{2}-9x^{2}t^{2},-y(3x^{3}+yzt))$
and
$\rho^{z}=(z(3t^{3}+xyz),9x^{2}t^{2}-y^{2}z^{2},0,z(3x^{3}+yzt)).$
We see that the syzygies $\omega^{y}$ and $\omega^{z}$ are primitive, since in each case the four coordinate polynomials have no common factor. Moreover, the syzygies of the form
$\rho=c^{y}\rho^{y}+c^{z}\rho^{z}+b\rho_{1}$
with $c^{y},c^{z}\in\mathbb{C}$ not both zero, and $b\in S_{3}$ are all primitive. Indeed, looking at the first and the fourth coordinate polynomials of $\rho$, we see that the only possible common factor is a linear form in $y,z$, but such a factor cannot divide the second and the third coordinates of $\rho$, due to the monomial $x^{2}t^{2}$ present there. It follows that one can take $\rho_{1}$, $\rho_{2}=\rho^{y}$ and $\rho_{3}=\rho^{z}$ as the first 3 generating syzygies for $AR(f)$ and hence $d_{1}=1$ and $d_{2}=d_{3}=4$ for any $m\geq 2$. This shows that
t(X)=1+4+4+13m=103mt(X)=1+4+4+1-3m=10-3m
can be an arbitrarily large negative integer. In this example the vector space spanned by the forms $\omega^{x},\omega^{y},\omega^{z}$ and $\omega^{t}$ has dimension 4, as predicted by Theorem 6.1. On the other hand, the obvious relation
$(9x^{2}t^{2}-y^{2}z^{2})\rho_{1}+z\rho_{2}+y\rho_{3}=0$
explain why Proposition 4.1 cannot be applied to these 3 syzygies and get $t(X)\geq 0$.